Operations Report – February 22nd

Crew 293 Operations Report 22-02-2024


Name of person filing report: Leo Tokaryev

Non-nominal systems: Suit n°2 and Suit n°6

Notes on non-nominal systems:

Suit n°2: one of two attachments that is supposed to hold the red circle that is holding the helmet is broken, we suppose that it broke during yesterday’s EVA.

Suit n°6: the suit has a serious problem with power. After 3 hours of EVA, the suit was on 9V, which made the ventilation system inefficient and caused a medical event right at the end of the EVA. The suit loses energy too fast, and still is too dangerous to use even though the battery was replaced by the previous crew.


Spirit rover used: No

Hours: (before EVA)

Beginning charge: (Before EVA)

Ending charge: (On return from EVA, before recharging)

Currently charging:

Opportunity rover used:Yes

Hours: 164.2

Beginning charge:100

Ending charge:72

Currently charging:Yes

Curiosity rover used: No


Beginning charge:

Ending charge:

Currently charging:

Perseverance rover used:Yes


Beginning charge:95

Ending charge:61

Currently charging:Yes

General notes on rovers:

Summary of Hab operations: put notes here

WATER USE: 9.8 gallons

Water (static tank): 407.9 gallons

Static tank pipe heater (on or off):On

Static tank heater (On or off)On

Toilet tank emptied: yes

Summary of internet: none

Summary of suits and radios: The radios were functioning well.

Summary of GreenHab operations: put notes here

WATER USE: 6 gallons

Heater: On

Supplemental light: 22h-2h on automatic

Harvest: nothing

Summary of ScienceDome operations: put notes here

Dual split: Heat or AC, On

Summary of RAM operations: RAM wasn’t used today

Summary of any observatory issues: none

Summary of health and safety issues: see the HSO report about the medical event that was caused by the malfunction of one of the EVA suits.

Questions, concerns and requests to Mission Support:

1) We decided to test all the suits’ battery usage, we will inform you after we have our results. Our protocol is to put all the suits on maximum ventilation mode and take voltage and air flow measurements every 30 minutes to test their battery usage. We will conduct this testing for 4 hours since it is the maximum time of an EVA. I believe that previous tests that were made on suit N°6 were conducted only for one hour, that is why the problem wasn’t detected. We will send you the full results, but for the moment we already identified some problems in suits 4 and 6.

We had an issue with toilets today: the odors were too strong not only in the restroom, but in the whole lower deck and even in some rooms in the upper deck. We emptied the toilets hoping that it would solve the problem, but the sensor only indicated that the tank was 1/3 full. I don’t know what is the source of the problem: the ventilation, the toilets or the sensor that wasn’t indicating the right value.

Copyright © The Mars Society. All rights reserved. | Main Site